1.) A 0.700 g sample of a nonvolatile yellow crystalline solid is dissolved in 16.3 g of benzene, producing a solution that freezes at 5.10oC. Find the molar mass of the yellow solid in g/mol.

2.) What is the osmotic pressure in atm of a 0.130 M solution of NaCl at 0oC?

3.) Calculate the freezing point for a 1.55 m solution of CCl4 in benzene.
Answer #1

Mass of solute = 0.7 g

Mass of solvent, benzene = 16.3 g = 0.0163 kg

Freezing point of pure benzene = 5.5 oC

Freezing point of solution = 5.1 oC

depression in freezing point = \Delta T_{f} = 5.5 - 5.1 = 0.40 oC

Kf of benzene = 4.90 oC/m

Formula

\Delta T_{f}=K_{f}\times molality

Solving for molality

Plug in the values

0.40 = 4.90 x m

molality = 0.082 m

But molality = moles/ weight of solvent

0.082 = moles/0.0163

multiply both sides with 0.0163

moles of solute = 0.082*0.0163 = 1.33 x 10-3 mol

Finding molar mass of solute

Molar mass of solute = weight /moles

= 0.7 g/1.33 x 10-3 mol

= 526.3 g/mol

---------------------------------------------------------------------------------------

2)

Concentration of NaCl solution = 0.130 M

Number of ions in NaCl = i = 2 (Na+ and Cl- ions)

R = 0.0821 L.atm/mol.K

Temperature = T = 0 oC = 273 K

Formula for osmotic pressure

\pi =i \times c\times R\times T

\pi =2\times0.130\times 0.0821\times 273

\pi =5.83 atm

------------------------------------------------------------------

3)

Given, molality of solution = 1.55 m

Kf = 4.90 oC/m

Formula for depression in freezing point =

\Delta T_{f}=K_{f}\times molality

\Delta T_{f}=4.90\times 1.55

\Delta T_{f}=7.60 ^{\circ}C

Freezing point of pure benzene = 5.5 oC

Freezing point of CCl4 -benzene solution = 5.50 - 7.60 = -2.1 oC